OpenStudy (anonymous):

At 10AM Eric traveled 195 miles across the plains states on his way to california. At 2 pm, he had traveled 455 miles. It is said that use slope to calculkate his rate of travel

OpenStudy (anonymous):

slope formula = change y / change x y is miles x is time create two ordered points (x1,y1) (x2,y2) change y = y2-y1 change x = x2-x1

OpenStudy (anonymous):

ah thanks for the formula